a) What is the momentum of a 10,000-kg truck whose speed is 30 m/s? b) What will be the speed of another 5,000-kg truck in order to have same momentum?


1
Expert's answer
2021-05-12T14:58:32-0400

(a)

"p=mv=10000\ kg\cdot30\ \dfrac{m}{s}=3\cdot10^5\ \dfrac{kgm}{s}."

(b)

"v=\dfrac{p}{m}=\dfrac{3\cdot10^5\ \dfrac{kgm}{s}}{5000\ kg}=60\ \dfrac{m}{s}."